what are the dimensions of a rectangle if the perimeter is 42 cm and the area is 80 cm squared?
Answer by Alan3354(69443)

L + W = 42/2 = 21 so one of the numbers has to be odd.
The only odd factors of 80 are 1 & 5.
1 by 20 doesn't work
--> 5 by 16